OpenStudy (anonymous):

Find the linearization of cos(2x) at pi/4. The answer was pi/2 - 2x but I got pi/4 -x?

OpenStudy (freckles):

what did you get when you evaluated cos(2x) at x=pi/4 and what did you get when you evaluated the derivative of cos(2x) at x=pi/4 ?

OpenStudy (freckles):

@amourette

OpenStudy (anonymous):

when i evaluated it the first time, I got 0, and when I derived it I got -1

OpenStudy (freckles):

cos(2x) evaluated at x=pi/4 is 0 so you are right there but -2sin(2x) evaluated at x=pi/4 is not -1...it is -2 I think you just forgot to multiply by 2...

OpenStudy (freckles):

\[y=f'(\frac{\pi}{4})(x-\frac{\pi}{4})+f(\frac{\pi}{4}) \\ \text{ where } f(x)=\cos(2x) \\ \text{ and } f'(x)=-2 \sin(2x) \text{ not just } -\sin(2x) \text{ which I what I believe you put maybe }\]

OpenStudy (anonymous):

why do you multiply it by 2?

OpenStudy (freckles):

oh maybe you don't know chain rule?

OpenStudy (anonymous):

ohhhh that's probably why. so when you're deriving -sin(2x), the 2 goes on the outside?

OpenStudy (freckles):

\[f(x)=\cos(2x) \\ f'(x)=(2x)' \cdot (\cos)'(2x) \\ f'(x)=2 \cdot [-\sin(2x)]\]

OpenStudy (freckles):

derivative of inside * derivative of outside

OpenStudy (anonymous):

I see my mistake thank you very much!

OpenStudy (freckles):

np